OpenStudy (anonymous):

is this true 7+x=84 but i think x=77

OpenStudy (anonymous):

7+x=84 x=84-7 x=77 you are correct

OpenStudy (anonymous):

Yes x=77. Once you subtract 7 on both sides, that leaves you with x=77 x+7=84 -7 -7 --------- x=77

OpenStudy (anonymous):

thank both of you

OpenStudy (anonymous):

No problem :D